OpenStudy (marinedude42):

A vertical line has a(n) ________ slope. A. undefined B. negative C. zero D. positive

OpenStudy (anonymous):

A

OpenStudy (marinedude42):

Find the slope of the line containing the points (-2, -2) and (-4, -2). A. -1 B. 1 C. 0 D. undefined

OpenStudy (mathstudent55):



OpenStudy (marinedude42):

??

OpenStudy (mathstudent55):

That was a vertical line. In a vertical line, all points have the same x-coordinate. When you calculate the slope, you can use any two points, but since you always divide by the difference of x values, the slope will involve division by zero, which is undefined. That's why the slope of a vertical line is undefined.

OpenStudy (mathstudent55):

For your second question, to find the slope, find the difference of the y-coordinates. Then find the difference of the x-coordinates. Divide the first difference by the second difference.

OpenStudy (marinedude42):

i got -2, and 0

OpenStudy (mathstudent55):

Be careful with the order. Do the difference of y's first: -2 - (-2) = -2 + 2 = 0 Now the difference of x's : -2 - (-4) = -2 + 4 = 2 Now divide the diff in y by the diff in x: slope = 0/2 = 0 slope = 0

OpenStudy (marinedude42):

i cant divide -2 by 0

OpenStudy (mathstudent55):

You don't need to divide -2 by 0. In this problem you need to divide 0 by -2, which you can do, and whose result is 0.

OpenStudy (marinedude42):

ok, thanks

OpenStudy (marinedude42):

In what direction is the line containing the point (5, 4) and (-2, 4) going? A. horizontal B. vertical C. up and to the right D. down and to the right

OpenStudy (mathstudent55):

Memorize this: Positive slope: up to the right Negative slope: down to the right Zero slope: horizontal Undefined slope: vertical

OpenStudy (marinedude42):

so up and to the right?

OpenStudy (mathstudent55):